### v1

# The Open Proof Corpus: A Large-Scale Study of LLM-Generated Mathematical Proofs

Abstract
In recent months, large language models (LLMs) have made significant progress in mathematical proof generation, but further advancement is hindered by the lack of a large-scale, high-quality dataset of human-evaluated proofs. While expensive to create, such a dataset is essential for driving improvements in training and enabling a rigorous analysis of proof generation capabilities. In this work, we present the Open Proof Corpus (OPC), a dataset comprising over 5,000 human-evaluated proofs produced by state-of-the-art LLMs. The OPC was specifically designed for broad applicability and downstream usage in proof generation research and is the first to include a substantial number of correct, LLM-generated solutions to problems from prestigious mathematics competitions such as the USAMO and IMO. Using the OPC, we explore critical questions in automated proof generation: (1) the performance gap between natural language and formal proof generation, (2) the discrepancy between final-answer accuracy and full-proof validity, and (3) the impact of best-of-n selection on proof quality. Finally, to showcase the utility of the OPC, we finetune an 8B-parameter model on the dataset, obtaining a model that performs on par with the best model, Gemini-2.5-Pro, on the task of evaluating proof correctness.
Machine Learning, ICML
Authors:
Dataset Contributors:
1 Introduction
Large language models (LLMs) have recently achieved remarkable progress in mathematical reasoning, attaining top-competitor performance on various final-answer benchmarks such as AIME and HMMT (Balunović et al., 2025). However, growing evidence suggests that these benchmarks fail to capture the full breadth of mathematical reasoning capabilities, as they do not require models to produce complete proofs or detailed intermediate steps (Mahdavi et al., 2025; Guo et al., 2025b). Such step-by-step reasoning is critical for applications in theorem proving, mathematical research, and education.
Proof benchmarking
To address this gap, several evaluation efforts have emerged, revealing that LLMs significantly underperform on proof generation compared to existing final-answer benchmarks (Petrov et al., 2025; Mahdavi et al., 2025; Guo et al., 2025b). Despite their value, these benchmarks are severely limited in their use for broader analysis, training, and future development purposes. Specifically, they are small in size (Petrov et al., 2025; Guo et al., 2025b), rely on outputs from outdated models (Frieder et al., 2023), contain few correct proofs (Mahdavi et al., 2025), and are not fully open-sourced (Mahdavi et al., 2025; Guo et al., 2025b).
Open questions
Furthermore, key questions about proof generation capabilities remain unanswered. First, while it is widely claimed that final-answer benchmarks are insufficient for evaluating proof generation capabilities, this claim is not yet substantiated by evaluating LLM-generated proofs on such benchmarks. Second, with recent advances in formal proof generation using systems like Lean (Ren et al., 2025; Wang et al., 2025), the performance gap between natural language and formal proof generation remains unclear. Third, the potential of best-of-n sampling strategies to improve proof quality has not been explored.
Sample Problem: EGMO 2012, Problem 3
Level: High School
Split: Generic
Question: Solve over $\mathbb{R}$ the functional equation $f(yf(x+y)+f(x))=4x+2yf(x+y)$ Refer to caption Proof: To solve the functional equation …we conclude that this is the **only** solution. Refer to caption Grade (Incorrect): The model substitutes $x$ to be 1. All further ideas are based on the wrong assumption. Refer to caption
((a)) Sample from the OPC
((b)) Conclusions from the OPC
Figure 1: Overview of the OPC and its conclusions. On the left, we show a typical sample from the OPC, including a question taken from a high-quality mathematical competition, a proof generated by an LLM, and a human judgment of the proof’s correctness. On the right, we summarize the main conclusions drawn from the OPC.
Our work: the Open Proof Corpus
To address these challenges, we introduce the Open Proof Corpus (OPC): a large-scale, human-validated dataset comprising over 5,0005{,}000 LLM-generated proofs across more than 1,0001{,}000 problems. As shown in Fig.˜1(a), each OPC sample includes (1) a problem from a high-quality mathematical competition such as the International Mathematical Olympiad (IMO), (2) a proof generated by a state-of-the-art LLM, and (3) a binary human evaluation of the proof’s correctness with feedback. The OPC was specifically designed for downstream usage in proof generation research, enabling both training and evaluation of LLMs on proof generation tasks. To address the open questions outlined earlier, the OPC includes problems from specific sources: problems from the PutnamBench (Tsoukalas et al., 2024) enable comparisons between formal and informal reasoning, while problems from MathArena (Balunović et al., 2025) support evaluation of proof correctness for problems with final answers.
Key findings
Despite the difficulty of problems in the OPC, state-of-the-art LLMs demonstrate surprisingly strong performance. For instance, o4-mini correctly solves almost $20\%$ of the problems in the IMO Shortlist, and the OPC generally consists of $43\%$ correct proofs. Furthermore, as shown in the top left of Fig.˜1(b), LLMs exhibit strong capabilities in evaluating proofs: Gemini-2.5-Pro achieves $88.1\%$ accuracy in judging proof correctness—nearly matching human performance at $90.4\%$. To showcase the utility of the OPC, we fine-tune DeepSeek-R1-Qwen3-8B (Guo et al., 2025a) using GRPO (Shao et al., 2024) on the OPC, resulting in an open model that achieves $88.1\%$ judgment accuracy, on-par with Gemini-2.5-Pro and outperforming all other models.
Answering open questions
The OPC allows us to empirically resolve the open questions outlined above, with all conclusions shown in Fig.˜1(b). First, as shown on the top right, we find that natural language proof generation significantly outperforms formal proof generation: on PutnamBench, Gemini-2.5-Pro solves 11 times more problems than the best formal model, Deepseek-Prover-V2 (Ren et al., 2025). Second, as shown on the bottom left, we observe a substantial gap between final-answer accuracy and proof correctness. While Gemini-2.5-Pro loses only $8\%$ of its final-answer accuracy when proof correctness is required, o3 suffers a drop of almost $30\%$. Third, as shown on the bottom right, we find that best-of-n strategies yield large gains in performance. Interestingly, while standard best-of-n selection methods moderately improve accuracy from $26\%$ to $36\%$, a ranking-based approach inspired by (Liu et al., 2025) achieves the highest performance of $47\%$.
Main Contributions
Our key contributions are:
  • A scalable pipeline for generating and evaluating natural language mathematical proofs (Section˜3).
  • The Open Proof Corpus, a large, human-validated dataset of over 5,000 LLM-generated proofs (Section˜4).
  • New benchmarking results on the OPC, resulting in novel insights into current LLM capabilities (Section˜5).
  • An open-source, 8B-parameter model fine-tuned on the OPC that achieves $88.1\%$ judgment accuracy, tied with the best model on this task (Section˜5).
2 Related Work
We briefly recapitulate the relevant literature on mathematical reasoning benchmarks and datasets.
Final-answer benchmarks
Final-answer benchmarks evaluate models by comparing a final answer from the model’s output with a ground-truth answer, typically using rule-based parsers. With the advent of reasoning LLMs (Jaech et al., 2024; Guo et al., 2025a), older benchmarks have become saturated (Cobbe et al., 2021; Lightman et al., 2024), and even more recent ones are nearing saturation (Gao et al., 2025; He et al., 2024; Balunović et al., 2025; Gulati et al., 2024). Only private benchmarks such as FrontierMath (Glazer et al., 2024) and HLE (Phan et al., 2025) remain sufficiently challenging for the latest models. However, these benchmarks do not require the generation of full proofs or detailed reasoning steps. Moreover, their private nature hinders reproducibility and broader community engagement, both key factors in driving progress.
Formal proof generation
Another growing line of work involves training LLMs to generate formal proofs in languages such as Lean (de Moura and Ullrich, 2021) or Isabelle (Nipkow et al., 2002), which can then be automatically verified by these systems (Zheng et al., 2022; Tsoukalas et al., 2024; Yu et al., 2025). While this paradigm enables rigorous correctness checking, it typically requires models to be specifically finetuned for formal proof generation (Ren et al., 2025; Wang et al., 2025; Lin et al., 2025). In contrast, state-of-the-art general-purpose models like o3 and Gemini-2.5-Pro struggle with formal proof generation. As a result, formal reasoning currently does not allow to make full use of the natural language capabilities of these general-purpose models, and, as we show in Section˜5, there remains a significant performance gap between formal and natural language proof generation, with the latter being substantially more effective.
Proof-generation evaluation efforts
Going beyond final-answer accuracy has gained recent attention, with several works investigating the reasoning traces of recent LLMs to identify patterns and potential for improvement (Shojaee et al., 2025; Mondorf and Plank, 2024; Xia et al., 2025). However, only a few studies have focused directly on evaluating full proofs. Petrov et al. (2025) evaluated LLMs on the six problems from the USAMO 2025, uncovering significant flaws in the generated proofs. Similarly, Mahdavi et al. (2025) evaluated LLM performance on a large set of IMO Shortlist problems, finding that no model surpassed $5\%$ accuracy. In contrast, Frieder et al. (2023) reported that GPT-3.5 and GPT-4 performed well on simpler tasks, generating correct proofs in a significant fraction of cases. Still, all these studies are limited by either the use of outdated models or the small scale of their evaluations.
Two concurrent works have recently expanded this line of research. Sheng et al. (2025) focus specifically on inequality proofs and primarily emphasize the development of an LLM as a judge framework to mitigate the high cost of human evaluation. Guo et al. (2025b) highlight a notable gap between final-answer accuracy and the ability to generate correct proofs, a finding we confirm in Section˜5. However, their analysis stops at this observation and does not evaluate performance on an established final-answer benchmark. Further, both studies are limited in scale compared to the OPC and have not open-sourced their human-annotated proofs.
Mathematical training datasets
Several large-scale datasets have been developed to train LLMs on mathematical reasoning. One of the earliest, Li et al. (2024), compiled a large dataset of internet-sourced problems, including both final-answer questions and natural language proofs. However, it lacks LLM-generated proofs, human judgments, and examples of incorrect proofs. Other datasets have focused exclusively on final-answer tasks (Albalak et al., 2025; He et al., 2025; Moshkov et al., 2025), offering limited support for training or evaluating proof generation. Finally, Zhang et al. (2025) introduced a dataset of both valid and invalid problem statements, each accompanied by LLM-generated proofs. While this ensures that incorrect proofs exist for invalid statements, the dataset does not include human evaluation of proofs or any other information on the validity of the proofs.
LLM as a judge
The widespread success of LLMs has led to the development of the "LLM as a judge" paradigm, which enables scalable, consistent, and cost-effective evaluation of complex outputs (Zheng et al., 2023; Gu et al., 2024). This approach serves as a promising alternative to traditional expert-based evaluation and opens up new possibilities for self-assessment during generation through methods such as best-of-n sampling, feedback-guided decoding, and self-training (Pan et al., 2023; Madaan et al., 2023; Weng et al., 2023; Jiang et al., 2025). However, concerns remain about the robustness of this framework in high-stakes domains like mathematical reasoning, where issues such as overconfidence and self-degradation can undermine reliability (Huang et al., 2024; Pan et al., 2023).
3 Methodology
Accurately evaluating LLM-generated proofs is a challenging task. Models often make difficult-to-detect errors, and they rarely acknowledge when they cannot solve a problem (see Section˜5). In this section, we outline the methodology used to create the OPC, with particular emphasis on the complexities of evaluating LLM-generated proofs and our efforts to maximize dataset size. Since human judges cannot reasonably spend hours studying each problem, we developed a pipeline to support efficient grading. This pipeline consists of three main components: problem and judge preparation (Section˜3.1), the grading procedure (Section˜3.2), and monitoring and validation (Section˜3.3).
3.1 Problem and Judge Preparation
Judge selection
Judges were selected from among former IMO participants or individuals who reached the final stages of IMO selection in their respective countries. We contacted each judge personally to ensure they were qualified, motivated, and easily accessible for ongoing communication. A total of $13$ judges were involved, each responsible for grading a varying number of problems. Three of the most active judges had prior experience with evaluating LLM-generated proofs. One judge served as the coordinator, maintaining regular communication, tracking progress, and ensuring consistency and motivation across the group.
Problem selection
Problems were drawn from top-tier national and international mathematics competitions, with the goal of capturing a balanced mix of correct and incorrect proofs. All problems were sourced from official materials. Non-English problems were translated using GPT-4.1 and manually verified for accuracy. When available, official ground-truth solutions were also extracted and provided to the judges as references.
Proof generation
Proofs were generated using a set of state-of-the-art LLMs known for their strength in mathematical reasoning. Specifically, we used o4-mini and o3 from OpenAI (OpenAI, 2025), Gemini-2.5-Pro from Google (Google DeepMind, 2025), Grok-3-Mini from xAI (xAI, 2025), Qwen3-235B-A22B from Qwen (Qwen Team, 2025), and the latest version of DeepSeek-R1 from DeepSeek (Guo et al., 2025a). DeepSeek-R1 was released mid-way through dataset construction and replaced Grok-3-Mini thereafter. We designed the model prompt to clearly instruct models to generate full solutions, refining it through small-scale pilot tests. The final prompt is shown in Section˜E.1. All models were run with default parameters and a maximum generation length of 64,00064{,}000 tokens. If this limit was exceeded before a complete proof was generated, we re-sampled. For the MathArena subset (Balunović et al., 2025), we only retained solutions with a correct final answer, retrying generation if necessary. For PutnamBench (Tsoukalas et al., 2024), we appended the informal final answer (if present) to the problem statement to mirror the setup for formal models, allowing direct comparison between natural and formal proof outputs.
3.2 Grading Procedure
User interface
We built a custom web interface to facilitate efficient grading. A sample instance is available at https://judgeproofs.xyz/sample, with screenshots included in App.˜D. The interface displays the problem, reference solution (if available), anonymized model-generated proof, and grading form. Judges could mark the problem or solution as malformed (to filter such cases from the dataset), grade the proof, and annotate sentences from the model-generated proof with comments. Continuous feedback from judges helped us refine the interface over time.
Judge instructions
Judges were asked to label proofs as either correct or incorrect and provide written justification. Precise grading guidelines were critical to avoid inconsistencies in edge cases, such as minor omissions or shortcuts. To prevent overly strict grading, we clearly defined what level of omissions and frequency of mistakes were acceptable in a correct proof. The full guidelines are available at https://judgeproofs.xyz/instructions/, with a summary in App.˜D. These instructions were developed collaboratively with the judges and finalized after a pilot phase.
Abstention and uncertainty
Judges were allowed to abstain from grading a proof if they lacked the necessary expertise or found the solution too complex or convoluted. They could also mark judgments as uncertain in borderline cases, which proved especially useful for near-correct proofs containing subtle errors. Less than $3\%$ of proofs in the OPC are flagged as uncertain.
LLM issue summaries
After several hundred graded proofs, we introduced a new feature to support grading: an automatically generated summary of the proof using o4-mini. These summaries flagged potential issues—such as logical gaps or missing steps—based on a specially designed prompt (see Section˜E.2). Importantly, the model was instructed not to give a final verdict, only to indicate possible issues for human review. Judges reported that this significantly improved their efficiency and accuracy in detecting errors. To ensure the inclusion of these summaries did not bias our judges, we evaluated the agreement rate between o4-mini as a judge and human graders before and after their introduction. There was no significant difference in agreement, suggesting no introduction of bias into the grading process. However, in experiments involving best-of-n selection—where the LLM judge acts as a selection mechanism—we omitted these summaries to avoid any form of compounding bias in the evaluation.
Problem distribution
Problems were assigned to judges based on their mathematical background and availability. Judges not qualified to evaluate undergraduate-level problems were excluded from grading those problems.
3.3 Monitoring and Validation
To ensure grading quality and judge consistency, we implemented a set of monitoring and validation procedures.
Coordinator
One experienced judge was assigned as coordinator, responsible for tracking grading progress, resolving issues, and ensuring motivation. As a senior author, the coordinator had detailed knowledge of the dataset’s goals and methodology and was available to answer judges’ questions throughout the process.
Pilot phase
Prior to full-scale grading, we conducted a test run with a limited number of problems. This helped us identify interface issues, unclear instructions, and other inefficiencies. Based on feedback, we revised the interface and grading guidelines before launching the full evaluation.
Monitoring discrepancies
Approximately $10\%$ of the proofs were double-graded to evaluate inter-judge consistency. Disagreements were reviewed by the coordinator to determine whether they arose from misunderstandings, ambiguous instructions, or errors overlooked by a judge. If possible, instructions were further improved to prevent similar issues. However, most inconsistencies stemmed from overlooked errors in the proofs and could therefore not be resolved by clarifying the instructions. If the coordinator identified a significant number of discrepancies for a specific judge, they would discuss the issue with the judge to clarify instructions.
4 Open Proof Corpus
Figure 2: Overview of competitions included in the OPC.
We now introduce the OPC, providing an overview of the competitions and models it includes and key dataset statistics. At a high level, the OPC is built to support the training and evaluation of proof generation models and to facilitate a rigorous analysis of their capabilities.
Basic properties
The OPC consists of 5,0625{,}062 proofs across 1,0101{,}010 distinct problems, generated by six state-of-the-art LLMs. Each proof is labeled as either correct or incorrect by one or two human judges. Labels are accompanied by short justifications, with optional annotations highlighting specific sentences within the proof. Each problem may also include metadata such as its competition source, difficulty level, and other relevant attributes.
Competitions
Problems were sourced from a wide range of prestigious mathematics competitions. A full breakdown is provided in App.˜A, with a summary shown in Fig.˜2. App.˜A also reports the average accuracy of the best-performing model per competition, offering a rough proxy for difficulty. Most problems are at the high school level, with only a small portion drawn from undergraduate-level competitions. Competitions were selected to cover a broad range of difficulties while ensuring an overall average model accuracy of approximately $50\%$.
Table 1: Number of solutions evaluated per model.
Model # Solutions
o4-mini 1615
o3 892
Qwen3-235B-A22B 890
Gemini-Pro 878
Grok-3-Mini 461
DeepSeek-R1 326
Models
Table˜1 summarizes the number of proofs generated by each model in the OPC. Not all models contributed to every problem, but most problems include solutions from at least five models. o4-mini contributed the largest share of proofs, as it was used extensively in best-of-n and pass@n experiments.
Human performance and noise
To estimate label reliability, we double-graded approximately $10\%$ of the dataset. Among these, judges agreed on the proof’s correctness in $90.4\%$ of cases. Assuming independent judgments, we can estimate the individual judge error rate $p$ by solving $0.904=(1-p)^{2}+p^{2}$, giving $p=5\%$. This indicates a relatively low noise level, which is expected for a human-annotated dataset of this complexity.
Data splits
The OPC is divided into four subsets, each serving a distinct purpose:
  • MathArena: A test set of $112$ problems from MathArena (Balunović et al., 2025), a final-answer benchmark for mathematical reasoning.
  • PutnamBench: A subset of $114$ problems from PutnamBench (Tsoukalas et al., 2024), used to compare natural language and formal proof generation.
  • Best-of-n: A subset of $152$ problems from the IMOSL, USAMO, and BMOSL, each solved multiple times by o4-mini. For 60 of these problems, all 8 generations were human-evaluated. The rest include judgments only for the generation selected by a best-of-n selection strategy.
  • Generic: A subset of $676$ problems from various competitions, including the IMOSL, USAMO, BMOSL, and others, all solved by multiple models.
The MathArena and PutnamBench subsets are drawn from existing public benchmarks and should be treated strictly as test sets. The generic and best-of-n subsets, by contrast, are intended for training, validation, and further analysis. However, a small portion of the generic subset is held back for benchmarking purposes as well.
5 Results
We present the main findings derived from the OPC. In Section˜5.1, we evaluate the proof-generation capabilities of various models. Section˜5.2 evaluates the ability of LLMs to judge the correctness of proofs. In Section˜5.3, we compare informal and formal proof-generation performance. Section˜5.4 analyzes proof correctness given correct final answers. Finally, Section˜5.5 examines the effectiveness of best-of-n selection strategies. In App.˜C, we additionally provide several qualitative observations about common mistakes in the generated proofs.
5.1 Gemini-2.5-Pro Is The Best Proof Generator
Figure 3: Average proof correctness of various models on the OPC. Data is split into two partitions, the first, resp. second, containing only problems answered by all models except DeepSeek-R1, resp. Grok-3-Mini. The discrepancy between the two partitions is due to the inclusion of harder competitions in the second partition.
Fig.˜3 shows the average proof correctness of each model on a filtered subset of the OPC, including only questions answered by all models except Grok-3-Mini or DeepSeek-R1. Gemini-2.5-Pro achieves the highest average accuracy, slightly outperforming o4-mini. In contrast, the two open-source models—Qwen3-235B-A22B and DeepSeek-R1—underperform significantly, highlighting the current performance gap between closed-source and open-source models in proof generation
Out of more than 1,700 incorrect solutions analyzed, models explicitly state their inability to solve the problem in only 114 instances, with all but five of those generated by o3. Even o3 is more likely to produce an incorrect proof than to acknowledge uncertainty. This widespread reluctance to admit failure highlights a key limitation: current LLMs lack effective mechanisms for confidence calibration and knowing their limitations. In domains like mathematics, this could undermine trust in systems that rely on LLMs for provably correct solutions, especially given the difficulty of evaluating proofs correctly.
5.2 LLMs Are Near-Human Level Judges
Table 2: Benchmarking LLMs as proof graders. Cost for running the model on the entire subset is given in USD.
Judge pass@1 maj@5 Cost
Human $90.4$ - N/A
Gemini-2.5-Pro $85.4$ $88.1$ $135.47$
OPC-R1-8B $83.8$ $88.1$ N/A
o4-mini $83.8$ $85.3$ $29.57$
o3 $83.1$ $84.3$ $93.3$
Gemini-2.5-Flash $82.7$ $86.0$ $86.95$
Qwen3-235B-A22B $81.8$ $84.6$ $3.79$
DeepSeek-R1 $80.9$ $82.6$ $27.35$
Qwen3-30B-A3B $74.0$ $75.4$ N/A
DeepSeek-R1-Qwen3-8B $70.7$ $71.3$ N/A
Claude-4-Sonnet $70.6$ $75.0$ $28.21$
Qwen3-8B $64.4$ $63.6$ N/A
GPT-4.1 $61.4$ $60.8$ $20.33$
Baseline $53.2$ $53.2$ N/A
The OPC is fundamentally a dataset of binary human judgments on proof correctness, making it well-suited for training and evaluating LLMs as proof judges. To leverage this, we split the generic subset by problem into training and test sets. Using GRPO (Shao et al., 2024), we fine-tune DeepSeek-R1-Qwen3-8B on the training set with human labels for reward modeling.
The test set comprises 293 LLM-generated proofs. We evaluate both reasoning models, such as Gemini-2.5-Pro and o4-mini, and general-purpose models like GPT-4.1. As a baseline, we classify proofs under 100 characters as incorrect and guess randomly otherwise.
As shown in Table˜2, Gemini-2.5-Pro achieves the highest judging accuracy: $85.4\%$ with a single evaluation pass and $88.1\%$ with majority voting, approaching the $90.4\%$ human baseline. Notably, OPC-R1-8B matches Gemini-2.5-Pro’s majority voting performance and outperforms its base model by $17\%$, demonstrating the value of the OPC and its potential for advancing the field of proof evaluation and generation. However, since the train and test sets share the same distribution, OPC-R1-8B’s performance may degrade on out-of-distribution data.
Table 3: Judgement accuracy breakdown split by solver, highlighting the lowest score for each judge. For visual clarity, we shortened Gemini-2.5-Pro to Gemini, o4-mini to o4, and Qwen3-235B-A22B to Qwen.
Judge Prover Gemini o4 o3 Qwen
Gemini-2.5-Pro 79.4 $87.1$ $91.6$ $80.6$
o4-mini $86.9$ 81.3 $83.1$ $84.1$
o3 $85.9$ $84.8$ 76.9 $87.8$
Qwen3-235B-A22B $80.0$ $81.9$ 79.1 $84.4$
LLMs often favor their own generations (Panickssery et al., 2024). To investigate this, we evaluate how well models judge their proofs compared to those generated by others. In Table˜3, we find that all models except Qwen3-235B-A22B perform significantly worse when judging their own proofs, indicating an inability to self-critique effectively. This suggests that LLMs struggle to recognize their mistakes, which is a critical limitation for applications requiring self-assessment or iterative improvement.
This positive result appears to contradict with Petrov et al. (2025), who reported poor performance of judge models. However, their evaluation was based on a limited set of problems, relied on older models, and focused on the more challenging task of scoring proofs on a continuous rather than a binary scale. In addition, we put considerable effort in crafting clear and comprehensible prompt instructions.
5.3 Formal Proof Generation Lags Behind
Figure 4: Average proof correctness on the PutnamBench.
Using the PutnamBench subset, we compare formal and natural language proof-generation models. The best formal model, DeepSeek-Prover-V2 (Ren et al., 2025), achieves less than $8\%$ accuracy on PutnamBench. In contrast, Fig.˜4 shows that the top informal model, Gemini-2.5-Pro, reaches almost $83\%$ accuracy on the evaluated subset, clearly outperforming all formal models. Despite this disparity, formal proofs offer a major advantage: automatic verifiability. While informal methods currently dominate in performance, formal approaches remain crucial for scalable, rigorous proof checking.
5.4 Proof Generation and Final Answer Do Not Align
Figure 5: Comparison of final-answer accuracy and proof correctness on the MathArena subset.
Although it is widely claimed that final-answer benchmarks are inadequate for evaluating proof generation capabilities (Petrov et al., 2025; Mahdavi et al., 2025; Guan et al., 2025), it remains unclear how often LLMs can produce a valid proof when they find the correct answer. To investigate this, we first collect instances from the MathArena subset where models generate correct final answers, and then manually evaluate the validity of the accompanying proofs. This enables us to estimate the overall proof correctness rate P(correct proof)P($\text${correct proof}) and compare it with the final-answer accuracy P(correct final answer)P($\text${correct final answer}).
In Fig.˜5, we report each model’s accuracy alongside the stricter metric requiring both a correct final answer and a valid proof. Despite Gemini-2.5-Pro, o4-mini, and o3 achieving similar final-answer accuracy, their proof correctness rates differ significantly. Specifically, o3 performs notably worse, with only $59.5\%$ of its answers containing a correct proof. This substantial variation across models highlights that final-answer accuracy alone is not a reliable indicator of proof generation capability.
5.5 Best-of-n Significantly Improves Performance
Figure 6: Pass@n metric on a small subset of the OPC compared with some best-of-n selection strategies.
Best-of-n selection—generating multiple outputs and selecting the best one—is a common strategy for improving LLM performance. We evaluate this approach using o4-mini by generating eight proofs per problem in the best-of-n subset and testing four selection methods:
  • Discrete: o4-mini classifies each proof as correct or incorrect and selects any correct proof.
  • Continuous: o4-mini scores proofs on a 0-77 scale and selects the one with the highest score.
  • Rank (Bracket): A ranking method proposed by Liu et al. (2025). Proofs are judged pairwise by o4-mini in a knockout tournament until one proof remains. Note that Liu et al. (2025) finetuned their model on the judging task, while we use o4-mini itself.
  • Rank (Swiss): Inspired by Liu et al. (2025), proofs are ranked via pairwise comparisons in a Swiss round-robin tournament. Ratings are computed using the Bradley-Terry model (Bradley and Terry, 1952), and the answer with the highest rating is selected. See App.˜B for details of this method.
Prompts for all methods can be found in App.˜E. Note that Rank (Bracket) only requires $O(n)$ comparisons, and is therefore as efficient as the discrete and continuous methods, while Rank (Swiss) requires $O(n^{2})$ comparisons, making it more expensive.
In Fig.˜6, we compare the performance of these methods with the pass@n metric on the $60$ problems that contain human judgments for all eight proofs. We find that best-of-n selection strategies can significantly improve proof generation performance. Furthermore, the pairwise ranking methods significantly outperform the discrete and continuous methods by approximately $10\%$. Notably, while discrete and continuous methods plateau after $n=5$, ranking approaches continue to scale. Finally, Rank (Swiss) slightly outperforms Rank (Bracket) by a $3\%$ margin.
Figure 7: Best-of-n selection strategies on the entire best-of-n subset of the OPC.
In Fig.˜7, we evaluate the performance of the best-of-n selection strategies on the entire best-of-n subset of the OPC, except for the Rank (Bracket) method, which was not evaluated on the full subset. The improved performance of the ranking methods is confirmed on the entire best-of-n subset, with Rank (Swiss) improving accuracy by $17\%$.
6 Limitations
While the OPC represents a significant advancement in the development and evaluation of LLM proof-generation capabilities, it is not without limitations. First, since most problems in the OPC are sourced from publicly available mathematical competitions, there is a risk of data contamination. Some of these problems may have been encountered during the training of the models. However, since the main purpose of the OPC is not related to benchmarking but rather to provide a large, human-validated dataset of LLM-generated proofs, this issue is less critical. Second, the majority of problems in the OPC are derived from high school level competitions. As a result, the dataset does not cover more advanced mathematical domains, such as undergraduate or research-level mathematics.
7 Future Work
The OPC offers a valuable resource for advancing research in proof generation with LLMs, supporting both the training and evaluation of models across a diverse set of mathematical problems. Our analysis also highlights several critical gaps and challenges that warrant further research. First, the significant disparity between formal and natural language proof generation—outlined in Section˜5.3—points to the need for more effective strategies to bridge this divide. Second, the OPC demonstrates that current benchmarks, which often rely solely on final answers, fail to capture the full complexity and quality of generated proofs, as discussed in Section˜5.4. This underscores the need for the development of a scalable benchmarking pipeline tailored to proof generation tasks. Finally, while our results show that best-of-n sampling strategies can meaningfully improve proof quality (see Section˜5.5), further research is required to better understand and optimize these methods.
8 Conclusion
In this work, we introduced the Open Proof Corpus (OPC), a large-scale, human-validated dataset of over 5,0005{,}000 LLM-generated proofs across more than 1,0001{,}000 problems. Designed to support training, evaluation, and benchmarking, the OPC provides a robust foundation for advancing research in automated proof generation. Using the OPC, we conducted a thorough evaluation of current LLM capabilities in mathematical reasoning. Our analysis reveals several key findings: (1) modern LLMs can generate a significant number of correct proofs; (2) some models approach human-level performance in judging proof correctness; (3) natural language proof generation substantially outperforms formal proof generation; (4) final-answer benchmarks are a poor proxy for proof quality; and (5) best-of-n sampling strategies significantly improve proof-generation accuracy. Using the OPC, we also trained an open-source model that achieves $88.1\%$ accuracy in judging proof correctness, matching the performance of the best closed-source model, Gemini-2.5-Pro.
Acknowledgements
This work has received funding from the Swiss National Science Foundation (SNSF) [200021_207967], the Ministry of Education and Science of Bulgaria (support for INSAIT, part of the Bulgarian National Roadmap for Research Infrastructure), and the Swiss State Secretariat for Education, Research and Innovation (SERI).
Appendix A Competitions in the OPC
The OPC contains over 1000 problems that were sourced from national and international competitions of varying difficulty. In Table˜4, we present the problem and sample distribution for each. We also include the following additional information:
  • Level: the education level the problems are appropriate for, either high school (HS) or undergraduate (UG).
  • Type: whether the competition is hosted internationally (I) or only nationally (N).
  • Source: we link the source, from which we obtained the problems. Any source that is not publicly available was marked as "Private".
  • Acc: the average accuracy of the best-performing model on the competition problems, which serves as a rough proxy for difficulty.
Table 4: A list of competition sources for the problems in OPC.
Competition Description Problems Solutions Level Type Acc Source
Main Analysis
Balkan MO Shortlist Competition between Balkan countries 74 368 HS I 31.7% Public
Baltic Way MO Northern and Central European Olympiad 80 395 HS I 68.1% Public
British MO Final Final round of the British Olympiad 23 114 HS N 78.3% Public
British MO Prelim Preliminary round for the British Olympiad 28 139 HS N 87.5% Public
Bulgarian Seasonal Competitions Seasonal Competitions hosted in Bulgaria (8th-12th grade) 49 242 HS N 63.5% Public
European Girls’ MO Europe-wide olympiad allowing only girls as participants 47 227 HS I 37.8% Public
IMC International competition for university students 43 212 UG I 62.2% Public
IMO Shortlist Shortlist of problems, from which the IMO is selected 128 629 HS I 18.0% Public
International Zhautykov MO Kazakhstan-based olympiad with near-IMO-level questions 41 203 HS I 34.2% Public
Irish MO Final round of the Irish Olympiad 51 255 HS N 92.0% Public
Putnam Undergraduate competition, regarded as one of the most difficult 8 35 UG I 100.0% Public
Romanian Masters of Mathematics Extralist IMO-level competition hosted in Romania 33 162 HS I 43.3% Private
Swiss MO Various problems from the Swiss Olympiad 8 39 HS N 87.5% Public
USA Junior MO USA olympiad that allows only junior students 25 121 HS N 52.2% Public
USAMO The final round of the USA Math Olympiad 38 190 HS N 35.1% Public
PutnamBench
Putnam Undergraduate competition, regarded as one of the most difficult 114 564 UG I 82.7% Public
MathArena
AIME 2025 Answer-based competition, serving as a qualifier for the USAMO 24 93 HS N 95.7% Public
BRUMO 2025 Answer-based competition hosted by Brown University 28 114 HS N 100.0% Public
HMMT February 2025 Answer-based competition hosted by Harvard and MIT 26 103 HS N 97.8% Public
SMT 2025 Answer-based competition hosted by Stanford 34 128 HS N 92.6% Private
Best-of-n
Balkan MO Shortlist Competition between Balkan countries 45 287 HS I 62.5% Public
IMO Shortlist Shortlist of problems, from which the IMO is selected 57 269 HS I 30.8% Public
USAMO The final round of the USA Math Olympiad 40 173 HS N 66.7% Public
Appendix B Swiss Ranking Methodology
We briefly describe the Swiss ranking method used as a best-of-n selection strategy. In this approach, a round-robin tournament is performed where each proof competes against every other. In each "game", two proofs are compared by o4-mini, which decides which proof is better, or if they are equally good.
To determine the overall winner, we compute a rating for each proof using the Bradley-Terry model [Bradley and Terry, 1952], a probabilistic model for paired comparisons commonly applied in LLM evaluation [Zheng et al., 2023, Dekoninck et al., 2025]. The Bradley-Terry model estimates the probability that a proof with rating $r_{i}$ beats a proof with rating $r_{j}$ as:
$P(i\text{ beats }j)=\frac{1}{1+\exp(r_{j}-r_{i})}.$
We fit the model to the outcomes of the round-robin tournament using maximum likelihood estimation, resulting in a rating for each proof. The proof with the highest rating is selected as the best.
Appendix C Qualitative Observations
Building on our quantitative analysis, we present several qualitative insights drawn from our judges’ evaluations. These include common failure modes and notable model behaviors.
Incorrect handling of inequalities
Many models sometimes make mistakes when manipulating inequalities. A frequent issue is the improper reversal of inequality signs during intermediate steps without mathematical justification. This results in incorrect final answers despite seemingly coherent reasoning. Below is an example of such a mistake made by DeepSeek-R1, where the incorrect sign reversal is marked with "**".
[Uncaptioned image]  DeepSeek-R1 – Baltic Way 2004 P3
Struggles with synthetic geometry reasoning
In Olympiad-level geometry problems, models often default to coordinate-based methods, i.e., using computations with complex numbers or a specific coordinate system to "bash" the problem, especially on easier problems where such techniques are tractable. However, for harder problems, models tend to shift toward synthetic approaches, relying on geometric theorems and properties. These harder problems typically require keeping track of multiple different but related angles and a robust understanding of geometric relationships. Here, models frequently struggle: they sometimes confuse angle orientations, misidentify congruent angles, reverse point orderings, or apply geometric properties incorrectly. In the following example, Gemini-2.5-Pro correctly identifies the base of an isosceles triangle in one case, but subsequently deduces the wrong pair of equal angles—illustrating a common angle-reasoning failure.
[Uncaptioned image]  Gemini-2.5-Pro – IMOSL 2013 G4
Incorrect citation of IMOSL 2004 G8 by Gemini-2.5-Pro
A curious behavior observed exclusively in Gemini-Pro is the repeated and inappropriate citation of the 2004 IMO Shortlist geometry problem G8. This reference appears in over 15 different solutions, often in problems unrelated to geometry. Notably, these citations consistently occur at points where the model appears to get stuck. No other model in our evaluation displayed this behavior. An illustrative example is shown below.
EGMO 2018 P3 [Uncaptioned image]  Gemini-2.5-Pro
Overgeneralization from simple cases
Consistent with prior work [Petrov et al., 2025, Mahdavi et al., 2025], we observe that models often overgeneralize patterns from small or simple examples. This tendency leads to incorrect solutions, particularly when the identified properties do not extend to the general case. An example illustrating this pattern overreach is included below.
[Uncaptioned image]  Grok 3 Mini – IMO SL 2000 G4
Appendix D Grading Interface and Instructions
This appendix outlines the grading interface and the accompanying instructions provided to judges. The full interface and documentation can be accessed online at https://judgeproofs.xyz/sample and https://judgeproofs.xyz/instructions.
Judge ID
Each judge received a unique identifier, which served as their login credential on https://judgeproofs.xyz. This ID was used to track grading progress while maintaining judge anonymity in the resulting dataset. To facilitate discussion and resolve ambiguities, a shared communication channel was created between all judges.
Grading interface
The grading interface was designed for clarity and ease of use. Figs.˜8, 9 and 10 illustrate its main components. The left panel contains a navigation bar for switching between problems and competitions. The right panel displays the problem statement and the ground-truth solution, along with options for flagging issues in either. Below, the generated solution is shown, accompanied by an automated summary and potential issues identified by an LLM judge. Judges can then evaluate the solution using a grading form that allows them to:
  • Indicate whether the solution is correct or incorrect
  • Provide a brief justification
  • Highlight specific parts of the solution relevant to their decision
  • Indicate uncertainty or abstain from grading
Instructions
Judges received a set of guidelines detailing how to use the interface and evaluate the correctness of solutions. Of particular importance were the criteria for determining whether a proof should be marked correct:
Instructions for judges on when a proof is correct These instructions were developed collaboratively with the judges and refined iteratively based on their feedback, ensuring consistent grading across different problems and evaluators.
Figure 8: Main grading interface. The left panel provides navigation across problems and competitions. The right panel displays the problem and ground-truth solution, with options to report issues.
Figure 9: Interface for evaluating a generated solution. The problem is repeated for reference. The generated solution appears on the left, and the LLM’s summary and identified issues on the right.
Figure 10: Grading form. Judges indicate correctness, provide a justification, highlight relevant content, and optionally express uncertainty or abstain.
Appendix E Prompts
In this section, we provide the prompts used for various tasks in the OPC. The prompts are designed to be clear and concise, guiding the LLMs through the proof generation process while ensuring that they understand the requirements for correctness and clarity. In Section˜E.1, we present the prompts used for generating proofs. In Section˜E.2, we provide the prompt used to generate the LLM summary to aid human graders in identifying potential issues in the proof. In Section˜E.3, we present the prompt used for LLMs to judge the correctness of a proof, used in Section˜5.2. In Sections˜E.4, E.5 and E.6, we provide all prompts used for the LLMs in best-of-n sampling, as described in Section˜5.5.
E.1 Proof Generation Prompt
The following prompt is used for problems with no final answer:
Prompt The following prompt is used for problems with a final answer: Prompt
E.2 Issues Interface Prompt
Prompt
E.3 LLM as Judge Prompt
Prompt
E.4 LLM as Discrete Judge Prompt
Prompt
E.5 LLM as Continuous Judge Prompt
Prompt
E.6 LLM as Rank Judge Prompt
Prompt

### v2

@@
 In recent months, large language models (LLMs) have made significant progress in mathematical proof generation, but further advancement is hindered by the lack of a large-scale, high-quality dataset of human-evaluated proofs.
-While expensive to create, such a dataset is essential for driving improvements in training and enabling a rigorous analysis of proof generation capabilities.
+While expensive to create, such a dataset is essential for driving improvements in training and addressing key open questions in the field of automated proof generation.
+Specifically, it remains unknown (1) how large the gap is between natural language and formal proof generation, (2) how final-answer accuracy relates to full proof correctness, and (3) how best-of-n selection strategies can affect proof quality.
 In this work, we present the Open Proof Corpus (OPC), a dataset comprising over 5,000 human-evaluated proofs produced by state-of-the-art LLMs.
-The OPC was specifically designed for broad applicability and downstream usage in proof generation research and is the first to include a substantial number of correct, LLM-generated solutions to problems from prestigious mathematics competitions such as the USAMO and IMO.
-Using the OPC, we explore critical questions in automated proof generation: (1) the performance gap between natural language and formal proof generation, (2) the discrepancy between final-answer accuracy and full-proof validity, and (3) the impact of best-of-n selection on proof quality.
-Finally, to showcase the utility of the OPC, we finetune an 8B-parameter model on the dataset, obtaining a model that performs on par with the best model, Gemini-2.5-Pro, on the task of evaluating proof correctness.
-Machine Learning, ICML
-Authors:
-Dataset Contributors:
+The OPC was specifically designed for broad applicability and downstream usage in proof generation research and is the first large dataset of LLM-generated solutions to problems from prestigious mathematics competitions such as the USAMO and IMO.
+Using the OPC, we address the open questions outlined above and provide new insights into LLMs’ strengths and limitations in mathematical reasoning.
+Finally, to showcase the utility of the OPC, we finetune an 8B-parameter model on the dataset, obtaining a model that matches Gemini-2.5-Pro, and performs close to the best model, GPT-5, on evaluating proof correctness.
 1 Introduction
-Large language models (LLMs) have recently achieved remarkable progress in mathematical reasoning, attaining top-competitor performance on various final-answer benchmarks such as AIME and HMMT (Balunović et al., 2025).
-However, growing evidence suggests that these benchmarks fail to capture the full breadth of mathematical reasoning capabilities, as they do not require models to produce complete proofs or detailed intermediate steps (Mahdavi et al., 2025; Guo et al., 2025b).
+Large language models (LLMs) have recently achieved impressive progress in mathematical reasoning, obtaining top-competitor performance on various final-answer benchmarks such as AIME and HMMT (Balunović et al., 2025).
+However, growing evidence suggests that these benchmarks fail to capture the full range of mathematical capabilities, as they do not require models to produce proofs or detailed intermediate steps (Mahdavi et al., 2025; Guo et al., 2025b).
 Such step-by-step reasoning is critical for applications in theorem proving, mathematical research, and education.
 Proof benchmarking
-To address this gap, several evaluation efforts have emerged, revealing that LLMs significantly underperform on proof generation compared to existing final-answer benchmarks (Petrov et al., 2025; Mahdavi et al., 2025; Guo et al., 2025b).
-Despite their value, these benchmarks are severely limited in their use for broader analysis, training, and future development purposes.
-Specifically, they are small in size (Petrov et al., 2025; Guo et al., 2025b), rely on outputs from outdated models (Frieder et al., 2023), contain few correct proofs (Mahdavi et al., 2025), and are not fully open-sourced (Mahdavi et al., 2025; Guo et al., 2025b).
+To address this, several evaluation efforts have taken place, revealing that LLMs significantly underperform on proof generation compared to existing final-answer benchmarks (Petrov et al., 2025; Mahdavi et al., 2025).
+Despite their value, these benchmarks are severely limited in their use for broader analysis and training.
+Specifically, they are small (Petrov et al., 2025; Guo et al., 2025b), use outdated models (Frieder et al., 2023), contain few correct proofs (Mahdavi et al., 2025), and are not open-sourced (Mahdavi et al., 2025; Guo et al., 2025b).
 Open questions
 Furthermore, key questions about proof generation capabilities remain unanswered.
-First, while it is widely claimed that final-answer benchmarks are insufficient for evaluating proof generation capabilities, this claim is not yet substantiated by evaluating LLM-generated proofs on such benchmarks.
-Second, with recent advances in formal proof generation using systems like Lean (Ren et al., 2025; Wang et al., 2025), the performance gap between natural language and formal proof generation remains unclear.
-Third, the potential of best-of-n sampling strategies to improve proof quality has not been explored.
-Sample Problem: EGMO 2012, Problem 3
-Level: High School
-Split: Generic
-Question: Solve over $\mathbb{R}$ the functional equation $f(yf(x+y)+f(x))=4x+2yf(x+y)$ Refer to caption Proof: To solve the functional equation …we conclude that this is the **only** solution.
+First, while it is widely claimed that there is a gap between final-answer performance and proof generation capabilities, this claim has not yet been supported by evaluating LLM-generated proofs on existing final-answer benchmarks.
+Second, despite recent advances in formal proof generation with Lean (Chen et al., 2025; Lin et al., 2025b), the performance gap between natural language and formal proof generation remains unclear.
+Third, the potential of best-of-n selection strategies to improve proof quality has not been explored.
+Sample Problem: EGMO 2012, Problem 3 Level: High School Split: Generic Question: Solve over $\mathbb{R}$ the functional equation $f(yf(x+y)+f(x))=4x+2yf(x+y)$ Refer to caption Proof: To solve the functional equation …we conclude that this is the **only** solution.
 Refer to caption Grade (Incorrect): The model substitutes $x$ to be 1.
@@
 Figure 1: Overview of the OPC and its conclusions.
-On the left, we show a typical sample from the OPC, including a question taken from a high-quality mathematical competition, a proof generated by an LLM, and a human judgment of the proof’s correctness.
-On the right, we summarize the main conclusions drawn from the OPC.
+On the left, we show a typical sample, including a question from a high-quality mathematical competition, an LLM-generated proof, and a human judgment of proof correctness.
+On the right, we summarize the main conclusions from the OPC.
 Our work: the Open Proof Corpus
 To address these challenges, we introduce the Open Proof Corpus (OPC): a large-scale, human-validated dataset comprising over 5,0005{,}000 LLM-generated proofs across more than 1,0001{,}000 problems.
-As shown in Fig.˜1(a), each OPC sample includes (1) a problem from a high-quality mathematical competition such as the International Mathematical Olympiad (IMO), (2) a proof generated by a state-of-the-art LLM, and (3) a binary human evaluation of the proof’s correctness with feedback.
+As shown in Fig.˜1(a), each OPC sample includes (1) a problem from a high-quality mathematical competition such as the International Mathematical Olympiad, (2) a proof generated by a state-of-the-art LLM, and (3) a binary human evaluation of the proof’s correctness with feedback.
 The OPC was specifically designed for downstream usage in proof generation research, enabling both training and evaluation of LLMs on proof generation tasks.
@@
 For instance, o4-mini correctly solves almost $20\%$ of the problems in the IMO Shortlist, and the OPC generally consists of $43\%$ correct proofs.
-Furthermore, as shown in the top left of Fig.˜1(b), LLMs exhibit strong capabilities in evaluating proofs: Gemini-2.5-Pro achieves $88.1\%$ accuracy in judging proof correctness—nearly matching human performance at $90.4\%$.
-To showcase the utility of the OPC, we fine-tune DeepSeek-R1-Qwen3-8B (Guo et al., 2025a) using GRPO (Shao et al., 2024) on the OPC, resulting in an open model that achieves $88.1\%$ judgment accuracy, on-par with Gemini-2.5-Pro and outperforming all other models.
+Furthermore, as shown in the top left of Fig.˜1(b), LLMs exhibit strong capabilities in evaluating proofs: GPT-5 achieves $90.8\%$ accuracy in judging proof correctness, which is on-par with human performance on this task.
+To showcase the utility of the OPC, we fine-tune R1-Qwen3-8B (Guo et al., 2025a) using GRPO (Shao et al., 2024) on the OPC, resulting in an open model that achieves $88.1\%$ judgment accuracy, close to top models like GPT-5 and outperforming the majority of frontier models.
 Answering open questions
 The OPC allows us to empirically resolve the open questions outlined above, with all conclusions shown in Fig.˜1(b).
-First, as shown on the top right, we find that natural language proof generation significantly outperforms formal proof generation: on PutnamBench, Gemini-2.5-Pro solves 11 times more problems than the best formal model, Deepseek-Prover-V2 (Ren et al., 2025).
+First, as shown on the top right, we find that natural language proof generation significantly outperforms formal proof generation: on PutnamBench, Gemini-2.5-Pro solves 4 times more problems than the best formal model, Goedel-Prover-V2 (Lin et al., 2025b).
 Second, as shown on the bottom left, we observe a substantial gap between final-answer accuracy and proof correctness.
@@
 Third, as shown on the bottom right, we find that best-of-n strategies yield large gains in performance.
-Interestingly, while standard best-of-n selection methods moderately improve accuracy from $26\%$ to $36\%$, a ranking-based approach inspired by (Liu et al., 2025) achieves the highest performance of $47\%$.
+Interestingly, while standard best-of-n selection methods moderately improve accuracy from $26\%$ to $36\%$, a ranking approach inspired by (Liu et al., 2025) achieves the highest performance of $47\%$.
 Main Contributions
 Our key contributions are:
-  • A scalable pipeline for generating and evaluating natural language mathematical proofs (Section˜3).
-  • The Open Proof Corpus, a large, human-validated dataset of over 5,000 LLM-generated proofs (Section˜4).
-  • New benchmarking results on the OPC, resulting in novel insights into current LLM capabilities (Section˜5).
-  • An open-source, 8B-parameter model fine-tuned on the OPC that achieves $88.1\%$ judgment accuracy, tied with the best model on this task (Section˜5).
+  • A rigorous pipeline for generating and evaluating natural language proofs (Section˜3).
+  • The OPC, a human-validated dataset comprising over 5,000 LLM-generated proofs (Section˜4).
+  • Resolution of open questions in natural language proof generation using the OPC (Section˜5).
+  • An open-source, 8B-parameter model fine-tuned on the OPC that achieves $88.1\%$ judgment accuracy, close to the best model on this task (Section˜5).
 2 Related Work
-We briefly recapitulate the relevant literature on mathematical reasoning benchmarks and datasets.
+We briefly discuss the relevant literature on mathematical reasoning benchmarks and datasets.
 Final-answer benchmarks
 Final-answer benchmarks evaluate models by comparing a final answer from the model’s output with a ground-truth answer, typically using rule-based parsers.
-With the advent of reasoning LLMs (Jaech et al., 2024; Guo et al., 2025a), older benchmarks have become saturated (Cobbe et al., 2021; Lightman et al., 2024), and even more recent ones are nearing saturation (Gao et al., 2025; He et al., 2024; Balunović et al., 2025; Gulati et al., 2024).
-Only private benchmarks such as FrontierMath (Glazer et al., 2024) and HLE (Phan et al., 2025) remain sufficiently challenging for the latest models.
-However, these benchmarks do not require the generation of full proofs or detailed reasoning steps.
-Moreover, their private nature hinders reproducibility and broader community engagement, both key factors in driving progress.
+With the rise of reasoning LLMs (Jaech et al., 2024), older benchmarks have become saturated (Cobbe et al., 2021; Lightman et al., 2024), and more recent ones are nearing saturation (Gao et al., 2025; He et al., 2024; Balunović et al., 2025).
+Only private benchmarks such as FrontierMath (Glazer et al., 2024) remain sufficiently challenging for the latest models.
+However, this benchmark does not require the generation of full proofs or detailed reasoning steps.
+Moreover, its private nature hinders reproducibility and broader community engagement, both key factors in driving progress.
 Formal proof generation
 Another growing line of work involves training LLMs to generate formal proofs in languages such as Lean (de Moura and Ullrich, 2021) or Isabelle (Nipkow et al., 2002), which can then be automatically verified by these systems (Zheng et al., 2022; Tsoukalas et al., 2024; Yu et al., 2025).
-While this paradigm enables rigorous correctness checking, it typically requires models to be specifically finetuned for formal proof generation (Ren et al., 2025; Wang et al., 2025; Lin et al., 2025).
-In contrast, state-of-the-art general-purpose models like o3 and Gemini-2.5-Pro struggle with formal proof generation.
-As a result, formal reasoning currently does not allow to make full use of the natural language capabilities of these general-purpose models, and, as we show in Section˜5, there remains a significant performance gap between formal and natural language proof generation, with the latter being substantially more effective.
+While this paradigm enables rigorous correctness checking, it typically requires models to be specifically finetuned for formal proof generation (Ren et al., 2025; Wang et al., 2025; Lin et al., 2025a).
+In contrast, state-of-the-art general-purpose models like GPT-5 and Gemini-2.5-Pro struggle with formal proof generation.
+As a result, formal reasoning currently does not make full use of the natural language capabilities of these general-purpose models, and, as we show in Section˜5, there remains a significant performance gap between formal and natural language proof generation, with the latter being substantially more effective.
 Proof-generation evaluation efforts
@@
 Still, all these studies are limited by either the use of outdated models or the small scale of their evaluations.
-Two concurrent works have recently expanded this line of research.
+Two recent works have expanded this line of research.
 Sheng et al.
@@
 Further, both studies are limited in scale compared to the OPC and have not open-sourced their human-annotated proofs.
+Automated proof grading with LLMs as judges
+The growing capabilities of LLMs have led to the “LLM-as-a-judge” paradigm, which enables cost-effective evaluation of complex model outputs (Zheng et al., 2023; Gu et al., 2024; Chevalier et al., 2024).
+Recent work has extended this approach to mathematical proof evaluation (Sheng et al., 2025; Zhao et al., 2024; 2025).
+These studies report promising results, showing that LLMs can grade specific types of problems with reasonable reliability (Sheng et al., 2025; Zhao et al., 2024) and can provide useful feedback in educational settings (Zhao et al., 2024; 2025).
+This work builds on these insights, building a dataset of human-evaluated proofs to further improve LLM judging capabilities, as detailed in Section˜5.
 Mathematical training datasets
@@
 (2025) introduced a dataset of both valid and invalid problem statements, each accompanied by LLM-generated proofs.
-While this ensures that incorrect proofs exist for invalid statements, the dataset does not include human evaluation of proofs or any other information on the validity of the proofs.
-LLM as a judge
-The widespread success of LLMs has led to the development of the "LLM as a judge" paradigm, which enables scalable, consistent, and cost-effective evaluation of complex outputs (Zheng et al., 2023; Gu et al., 2024).
-This approach serves as a promising alternative to traditional expert-based evaluation and opens up new possibilities for self-assessment during generation through methods such as best-of-n sampling, feedback-guided decoding, and self-training (Pan et al., 2023; Madaan et al., 2023; Weng et al., 2023; Jiang et al., 2025).
-However, concerns remain about the robustness of this framework in high-stakes domains like mathematical reasoning, where issues such as overconfidence and self-degradation can undermine reliability (Huang et al., 2024; Pan et al., 2023).
+While this ensures that incorrect proofs exist for invalid statements, the dataset does not include human evaluation of proofs or other information on proof validity.
 3 Methodology
@@
 Since human judges cannot reasonably spend hours studying each problem, we developed a pipeline to support efficient grading.
-This pipeline consists of three main components: problem and judge preparation (Section˜3.1), the grading procedure (Section˜3.2), and monitoring and validation (Section˜3.3).
+This pipeline consisted of three main components: problem and judge preparation (Section˜3.1), the grading procedure (Section˜3.2), and monitoring and validation (Section˜3.3).
+The dataset was constructed over a period of four weeks, involving $13$ expert judges and generating over 5,0005{,}000 human-validated proofs.
 3.1 Problem and Judge Preparation
 Judge selection
-Judges were selected from among former IMO participants or individuals who reached the final stages of IMO selection in their respective countries.
-We contacted each judge personally to ensure they were qualified, motivated, and easily accessible for ongoing communication.
+Judges were selected from among former IMO participants or individuals who reached the final stages of IMO selection in their countries.
+We contacted each judge personally to ensure they were qualified and motivated.
 A total of $13$ judges were involved, each responsible for grading a varying number of problems.
@@
 Problem selection
-Problems were drawn from top-tier national and international mathematics competitions, with the goal of capturing a balanced mix of correct and incorrect proofs.
+Problems were drawn from top-tier national and international mathematics competitions, such as the USAMO and IMO, to capture a balanced mix of correct and incorrect proofs.
+In particular, we included competitions based on two informal criteria: (1) the competition is well-known and produces high-quality problems, and (2) the difficulty level of the competition aligns with our target of roughly 50% model accuracy.
 All problems were sourced from official materials.
-Non-English problems were translated using GPT-4.1 and manually verified for accuracy.
+Non-English problems were translated using GPT-4.1 and manually verified for accuracy by the coordinator.
 When available, official ground-truth solutions were also extracted and provided to the judges as references.
+Throughout the annotation process, model performance was actively monitored to ensure that the selected problems remained appropriately challenging.
+For instance, more problems from international competitions were added when initial results indicated that models were performing very well ($\approx 65\%$) on national-level problems.
+Each day, problem prioritization was adjusted based on ongoing performance metrics, judge availability, and progress towards the specific conclusions we aimed to draw from the dataset.
 Proof generation
 Proofs were generated using a set of state-of-the-art LLMs known for their strength in mathematical reasoning.
-Specifically, we used o4-mini and o3 from OpenAI (OpenAI, 2025), Gemini-2.5-Pro from Google (Google DeepMind, 2025), Grok-3-Mini from xAI (xAI, 2025), Qwen3-235B-A22B from Qwen (Qwen Team, 2025), and the latest version of DeepSeek-R1 from DeepSeek (Guo et al., 2025a).
-DeepSeek-R1 was released mid-way through dataset construction and replaced Grok-3-Mini thereafter.
+Specifically, we used o4-mini and o3 from OpenAI (OpenAI, 2025), Gemini-2.5-Pro from Google (Team, 2025), Grok-3-Mini from xAI (xAI, 2025), Qwen3-235B-A22B from Qwen (Qwen Team, 2025), and the latest version of R1 from DeepSeek (Guo et al., 2025a).
+R1 was released mid-way through dataset construction and replaced Grok-3-Mini thereafter.
 We designed the model prompt to clearly instruct models to generate full solutions, refining it through small-scale pilot tests.
-The final prompt is shown in Section˜E.1.
+The final prompt is shown in Section˜I.1.
 All models were run with default parameters and a maximum generation length of 64,00064{,}000 tokens.
-If this limit was exceeded before a complete proof was generated, we re-sampled.
 For the MathArena subset (Balunović et al., 2025), we only retained solutions with a correct final answer, retrying generation if necessary.
@@
 We built a custom web interface to facilitate efficient grading.
-A sample instance is available at https://judgeproofs.xyz/sample, with screenshots included in App.˜D.
+A sample instance is available in our supplementary material, with screenshots included in App.˜H.
 The interface displays the problem, reference solution (if available), anonymized model-generated proof, and grading form.
@@
 To prevent overly strict grading, we clearly defined what level of omissions and frequency of mistakes were acceptable in a correct proof.
-The full guidelines are available at https://judgeproofs.xyz/instructions/, with a summary in App.˜D.
-These instructions were developed collaboratively with the judges and finalized after a pilot phase.
+The full guidelines are available in our supplementary material, with a summary in App.˜H.
+These instructions were developed collaboratively with the judges and finalized after a pilot phase (see Section˜3.3).
 Abstention and uncertainty
@@
 After several hundred graded proofs, we introduced a new feature to support grading: an automatically generated summary of the proof using o4-mini.
-These summaries flagged potential issues—such as logical gaps or missing steps—based on a specially designed prompt (see Section˜E.2).
-Importantly, the model was instructed not to give a final verdict, only to indicate possible issues for human review.
+These summaries flagged potential issues, such as logical gaps or missing steps, based on a specially designed prompt (see Section˜I.2).
+Importantly, the model was instructed not to give a final verdict.
 Judges reported that this significantly improved their efficiency and accuracy in detecting errors.
 To ensure the inclusion of these summaries did not bias our judges, we evaluated the agreement rate between o4-mini as a judge and human graders before and after their introduction.
-There was no significant difference in agreement, suggesting no introduction of bias into the grading process.
-However, in experiments involving best-of-n selection—where the LLM judge acts as a selection mechanism—we omitted these summaries to avoid any form of compounding bias in the evaluation.
+There was no significant difference in agreement, indicating that no bias was introduced into the grading process.
+However, in experiments involving best-of-n selection, where the LLM judge acts as a selection mechanism, we omitted these summaries to avoid any form of compounding bias in the evaluation.
 Problem distribution
-Problems were assigned to judges based on their mathematical background and availability. Judges not qualified to evaluate undergraduate-level problems were excluded from grading those problems.
+Problems were assigned to judges based on background and availability.
+Judges not qualified to evaluate undergraduate-level problems were excluded from grading them.
+To ensure a balanced workload, we monitored grading progress and dynamically reassigned problems as needed.
 3.3 Monitoring and Validation
-To ensure grading quality and judge consistency, we implemented a set of monitoring and validation procedures.
+To ensure grading quality and judge consistency, we implemented a set of validation procedures.
 Coordinator
 One experienced judge was assigned as coordinator, responsible for tracking grading progress, resolving issues, and ensuring motivation.
-As a senior author, the coordinator had detailed knowledge of the dataset’s goals and methodology and was available to answer judges’ questions throughout the process.
+As a core author, the coordinator had detailed knowledge of the dataset’s goals and methodology and was available to answer judges’ questions.
+Developer
+In addition to the coordinator, one of the core authors served as developer, providing technical support for the grading interface and addressing any software-related issues that arose during the grading process.
+Judges were encouraged to report bugs or suggest improvements to the grading interface, which the developer would address immediately to ensure a smooth grading experience.
+The developer was also responsible for implementing new features, such as the LLM-generated issue summaries, and to monitor model performance metrics to inform problem selection and judge assignments.
 Pilot phase
-Prior to full-scale grading, we conducted a test run with a limited number of problems.
-This helped us identify interface issues, unclear instructions, and other inefficiencies.
-Based on feedback, we revised the interface and grading guidelines before launching the full evaluation.
+Before full-scale grading, we conducted a test run with a limited number of problems.
+In particular, four experienced judges graded approximately 300 proofs in this initial stage.
+During this phase, a more significant portion of proofs (around $35\%$) were double-graded to evaluate inter-judge consistency.
+Judges were also very actively monitored, and encouraged to actively ask questions about any small ambiguity they encountered.
+After this stage, judge feedback was collected to improve the grading interface and instructions.
+In particular, the instructions were refined to address remaining ambiguities and edge cases.
 Monitoring discrepancies
-Approximately $10\%$ of the proofs were double-graded to evaluate inter-judge consistency.
-Disagreements were reviewed by the coordinator to determine whether they arose from misunderstandings, ambiguous instructions, or errors overlooked by a judge.
+In total, approximately $10\%$ of the proofs were double-graded to evaluate inter-judge consistency.
+Throughout the grading pipeline, disagreements were reviewed by the coordinator to determine whether they arose from misunderstandings, ambiguous instructions, or errors overlooked by a judge.
 If possible, instructions were further improved to prevent similar issues.
-However, most inconsistencies stemmed from overlooked errors in the proofs and could therefore not be resolved by clarifying the instructions.
+However, most inconsistencies came from overlooked errors in the proofs and could therefore not be resolved by clarifying the instructions.
 If the coordinator identified a significant number of discrepancies for a specific judge, they would discuss the issue with the judge to clarify instructions.
 4 Open Proof Corpus
-Figure 2: Overview of competitions included in the OPC.
-We now introduce the OPC, providing an overview of the competitions and models it includes and key dataset statistics.
-At a high level, the OPC is built to support the training and evaluation of proof generation models and to facilitate a rigorous analysis of their capabilities.
+We now introduce the OPC and provide an overview of key dataset statistics.
+At a high level, the OPC is built to support training and to facilitate a rigorous analysis of proof generation capabilities.
+Figure 2: Overview of competitions in the OPC.
 Basic properties
@@
 Each problem may also include metadata such as its competition source, difficulty level, and other relevant attributes.
+An example is shown in Fig.˜1(a).
 Competitions
@@
 App.˜A also reports the average accuracy of the best-performing model per competition, offering a rough proxy for difficulty.
-Most problems are at the high school level, with only a small portion drawn from undergraduate-level competitions.
-Competitions were selected to cover a broad range of difficulties while ensuring an overall average model accuracy of approximately $50\%$.
-Table 1: Number of solutions evaluated per model.
+Most problems are at the high school level ($\approx 84\%$), with a small portion drawn from undergraduate-level competitions ($\approx 16\%$).
+Table 1: Solutions per evaluated model.
 Model # Solutions
@@
 Models
-Table˜1 summarizes the number of proofs generated by each model in the OPC.
-Not all models contributed to every problem, but most problems include solutions from at least five models.
-o4-mini contributed the largest share of proofs, as it was used extensively in best-of-n and pass@n experiments.
-Human performance and noise
+Table˜1 shows the number of proofs generated by each model.
+Not all models attempted every problem, but most problems include solutions from five models.
+o4-mini contributed the largest number of proofs, as it was used in the best-of-n and pass@n experiments.
+Human performance
 To estimate label reliability, we double-graded approximately $10\%$ of the dataset.
@@
 Assuming independent judgments, we can estimate the individual judge error rate $p$ by solving $0.904=(1-p)^{2}+p^{2}$, giving $p=5\%$.
-This indicates a relatively low noise level, which is expected for a human-annotated dataset of this complexity.
+This indicates a low noise level, which is expected for a human-annotated dataset of this complexity.
 Data splits
 The OPC is divided into four subsets, each serving a distinct purpose:
-  • MathArena: A test set of $112$ problems from MathArena (Balunović et al., 2025), a final-answer benchmark for mathematical reasoning.
+  • MathArena: A subset of $112$ problems from MathArena (Balunović et al., 2025), a final-answer benchmark for mathematical reasoning.
+$34$ questions from the SMT 2025 will only be released after the competition authors have published the questions publicly.
   • PutnamBench: A subset of $114$ problems from PutnamBench (Tsoukalas et al., 2024), used to compare natural language and formal proof generation.
-  • Best-of-n: A subset of $152$ problems from the IMOSL, USAMO, and BMOSL, each solved multiple times by o4-mini.
+  • Best-of-n: A subset of $152$ problems from hard competitions, each solved multiple times by o4-mini.
 For 60 of these problems, all 8 generations were human-evaluated.
-The rest include judgments only for the generation selected by a best-of-n selection strategy.
-  • Generic: A subset of $676$ problems from various competitions, including the IMOSL, USAMO, BMOSL, and others, all solved by multiple models.
-The MathArena and PutnamBench subsets are drawn from existing public benchmarks and should be treated strictly as test sets.
-The generic and best-of-n subsets, by contrast, are intended for training, validation, and further analysis.
-However, a small portion of the generic subset is held back for benchmarking purposes as well.
+The rest include judgments only for the generations selected by a best-of-n selection strategy.
+  • Generic: A subset of $676$ problems from various competitions, solved by multiple models.
+The MathArena and PutnamBench subsets are drawn from existing benchmarks and should not be used for training.
+The generic and best-of-n subsets are intended for training, validation, and further analysis.
+However, a small portion of the generic subset is held back for benchmarking purposes.
 5 Results
-We present the main findings derived from the OPC.
+We now present our main findings.
 In Section˜5.1, we evaluate the proof-generation capabilities of various models.
-Section˜5.2 evaluates the ability of LLMs to judge the correctness of proofs.
+Section˜5.2 evaluates the ability of LLMs to judge proof correctness.
 In Section˜5.3, we compare informal and formal proof-generation performance.
 Section˜5.4 analyzes proof correctness given correct final answers.
-Finally, Section˜5.5 examines the effectiveness of best-of-n selection strategies.
-In App.˜C, we additionally provide several qualitative observations about common mistakes in the generated proofs.
-5.1 Gemini-2.5-Pro Is The Best Proof Generator
-Figure 3: Average proof correctness of various models on the OPC.
+Section˜5.5 examines the effectiveness of various best-of-n strategies.
+Finally, in Section˜5.6, we discuss the potential impact of data contamination on our results.
+In App.˜E, we additionally provide some qualitative observations about common mistakes in the generated proofs.
+All shown confidence intervals in this section are 95% intervals computed using the large sample normal approximation.
+Bold numbers indicate the best performance in the respective category, and underlined numbers are within the confidence interval of the best performance.
+5.1 Proof Generation Capabilities
+Figure 3: Average proof correctness on the OPC.
 Data is split into two partitions, the first, resp.
-second, containing only problems answered by all models except DeepSeek-R1, resp.
+second, containing only problems answered by all models except R1, resp.
 Grok-3-Mini.
-The discrepancy between the two partitions is due to the inclusion of harder competitions in the second partition.
-Fig.˜3 shows the average proof correctness of each model on a filtered subset of the OPC, including only questions answered by all models except Grok-3-Mini or DeepSeek-R1.
-Gemini-2.5-Pro achieves the highest average accuracy, slightly outperforming o4-mini.
-In contrast, the two open-source models—Qwen3-235B-A22B and DeepSeek-R1—underperform significantly, highlighting the current performance gap between closed-source and open-source models in proof generation
+The second partition contains problems from more challenging competitions, explaining the score discrepancy.
+Proof generation results
+Fig.˜3 shows the average proof correctness of each model on a subset of the OPC, including only fully-judged questions for all models except Grok-3-Mini or R1.
+Gemini-2.5-Pro achieves the highest accuracy, slightly outperforming o4-mini.
+In contrast, the two open-source models, Qwen3-235B-A22B and R1, underperform significantly, highlighting the performance gap between closed- and open-source models in generating correct proofs.
+Uncertainty acknowledgment
 Out of more than 1,700 incorrect solutions analyzed, models explicitly state their inability to solve the problem in only 114 instances, with all but five of those generated by o3.
 Even o3 is more likely to produce an incorrect proof than to acknowledge uncertainty.
-This widespread reluctance to admit failure highlights a key limitation: current LLMs lack effective mechanisms for confidence calibration and knowing their limitations.
-In domains like mathematics, this could undermine trust in systems that rely on LLMs for provably correct solutions, especially given the difficulty of evaluating proofs correctly.
-5.2 LLMs Are Near-Human Level Judges
-Table 2: Benchmarking LLMs as proof graders. Cost for running the model on the entire subset is given in USD.
+This reluctance to admit failure highlights a key limitation.
+In domains like mathematics, this could undermine trust in systems that rely on LLMs for provably correct solutions.
+5.2 LLMs Are Human Level Judges
+Setup
+The OPC provides binary human judgments for proof correctness, making it ideal for training and evaluating LLM proof judges.
+To leverage this, we split the generic subset by problem into training and test sets.
+Using GRPO (Shao et al., 2024), we fine-tune R1-Qwen3-8B using human labels for rewards.
+The test set comprises 293 LLM-generated proofs with 40% average correctness.
+We evaluate both reasoning models, such as GPT-5, and general-purpose models like GPT-4.1.
+Importantly, the human baseline is not measured on the test subset, but rather on all double-graded proofs in the OPC.
+Since the test samples are uniformly drawn from the OPC, this does not significantly affect the comparison.
+Table 2: LLMs as proof graders. Cost for running the model on the entire subset is given in USD. Full confidence intervals are given in Section˜C.4.
 Judge pass@1 maj@5 Cost
 Human $90.4$ - N/A
+GPT-5 89.3 90.8 $117.77$
+Grok-4 88.3 89.8 $104.42$
 Gemini-2.5-Pro $85.4$ 88.1 $135.47$
@@
 Qwen3-235B-A22B $81.8$ $84.6$ $3.79$
-DeepSeek-R1 $80.9$ $82.6$ $27.35$
-Qwen3-30B-A3B $74.0$ $75.4$ N/A
-DeepSeek-R1-Qwen3-8B $70.7$ $71.3$ N/A
+R1 $80.9$ $82.6$ $27.70$
+R1-Qwen3-8B $70.7$ $71.3$ N/A
 Claude-4-Sonnet $70.6$ $75.0$ $28.21$
@@
 GPT-4.1 $61.4$ $60.8$ $20.33$
-Baseline $53.2$ $53.2$ N/A
-The OPC is fundamentally a dataset of binary human judgments on proof correctness, making it well-suited for training and evaluating LLMs as proof judges.
-To leverage this, we split the generic subset by problem into training and test sets.
-Using GRPO (Shao et al., 2024), we fine-tune DeepSeek-R1-Qwen3-8B on the training set with human labels for reward modeling.
-The test set comprises 293 LLM-generated proofs.
-We evaluate both reasoning models, such as Gemini-2.5-Pro and o4-mini, and general-purpose models like GPT-4.1.
-As a baseline, we classify proofs under 100 characters as incorrect and guess randomly otherwise.
-As shown in Table˜2, Gemini-2.5-Pro achieves the highest judging accuracy: $85.4\%$ with a single evaluation pass and $88.1\%$ with majority voting, approaching the $90.4\%$ human baseline.
+Results
+As shown in Table˜2, GPT-5 achieves the highest judging accuracy: $89.3\%$ with a single evaluation pass, approaching the $90.4\%$ human baseline, and $90.8\%$ with majority voting.
 Notably, OPC-R1-8B matches Gemini-2.5-Pro’s majority voting performance and outperforms its base model by $17\%$, demonstrating the value of the OPC and its potential for advancing the field of proof evaluation and generation.
-However, since the train and test sets share the same distribution, OPC-R1-8B’s performance may degrade on out-of-distribution data.
-Table 3: Judgement accuracy breakdown split by solver, highlighting the lowest score for each judge.
-For visual clarity, we shortened Gemini-2.5-Pro to Gemini, o4-mini to o4, and Qwen3-235B-A22B to Qwen.
-Judge Prover Gemini o4 o3 Qwen
-Gemini-2.5-Pro 79.4 $87.1$ $91.6$ $80.6$
-o4-mini $86.9$ 81.3 $83.1$ $84.1$
-o3 $85.9$ $84.8$ 76.9 $87.8$
-Qwen3-235B-A22B $80.0$ $81.9$ 79.1 $84.4$
+However, the train set for OPC-R1-8B shares the same distribution as this test set, which may inflate its performance.
+In App.˜C, we show that while the performance of OPC-R1-8B is reduced under out-of-distribution data, the improvement over the base model persists even under these conditions.
+This positive result appears to contradict Petrov et al.
+(2025), who reported poor performance of judge models.
+However, their evaluation was based on a limited set of problems, relied on older models, and focused on the more challenging task of scoring proofs on a continuous rather than a binary scale.
+In addition, we put considerable effort into crafting clear and comprehensible prompts.
+Table 3: Judgement accuracy breakdown, highlighting the lowest score for each judge. Full confidence intervals are given in Section˜C.4.
+Prover Judge Gemini o4 o3 Qwen
+Gemini 79.4 $86.9$ $85.9$ 80.0
+o4 $87.1$ 81.3 $84.8$ 81.9
+o3 $91.6$ 83.1 76.9 79.1
+Qwen 80.6 84.1 $87.8$ $84.4$
+Self-evaluation
 LLMs often favor their own generations (Panickssery et al., 2024).
 To investigate this, we evaluate how well models judge their proofs compared to those generated by others.
-In Table˜3, we find that all models except Qwen3-235B-A22B perform significantly worse when judging their own proofs, indicating an inability to self-critique effectively.
-This suggests that LLMs struggle to recognize their mistakes, which is a critical limitation for applications requiring self-assessment or iterative improvement.
-This positive result appears to contradict with Petrov et al.
-(2025), who reported poor performance of judge models.
-However, their evaluation was based on a limited set of problems, relied on older models, and focused on the more challenging task of scoring proofs on a continuous rather than a binary scale.
-In addition, we put considerable effort in crafting clear and comprehensible prompt instructions.
+In Table˜3, we find that all models except Qwen3-235B-A22B perform worse when judging their own proofs.
+This suggests that LLMs struggle to recognize their own mistakes, which is a critical limitation for applications requiring self-evaluation.
 5.3 Formal Proof Generation Lags Behind
 Figure 4: Average proof correctness on the PutnamBench.
+Figure 5: Comparison of final-answer accuracy and proof correctness on the MathArena subset.
 Using the PutnamBench subset, we compare formal and natural language proof-generation models.
-The best formal model, DeepSeek-Prover-V2 (Ren et al., 2025), achieves less than $8\%$ accuracy on PutnamBench.
-In contrast, Fig.˜4 shows that the top informal model, Gemini-2.5-Pro, reaches almost $83\%$ accuracy on the evaluated subset, clearly outperforming all formal models.
+The best formal model, Goedel-Prover-v2 (Lin et al., 2025b), achieves less than $19\%$ accuracy on the PutnamBench.
+In contrast, Fig.˜5 shows that the top informal model, Gemini-2.5-Pro, reaches almost $83\%$ accuracy on the evaluated subset, clearly outperforming this baseline.
 Despite this disparity, formal proofs offer a major advantage: automatic verifiability.
-While informal methods currently dominate in performance, formal approaches remain crucial for scalable, rigorous proof checking.
+While informal methods currently dominate, formal approaches remain crucial for scalable proof verification.
+Finally, we note that a recent private agentic system, Seed-Prover (Chen et al., 2025), obtained a $50\%$ formal accuracy on the PutnamBench.
+However, our informal results do not use agentic techniques, and it is therefore not accurate to directly compare these numbers.
 5.4 Proof Generation and Final Answer Do Not Align
-Figure 5: Comparison of final-answer accuracy and proof correctness on the MathArena subset.
 Although it is widely claimed that final-answer benchmarks are inadequate for evaluating proof generation capabilities (Petrov et al., 2025; Mahdavi et al., 2025; Guan et al., 2025), it remains unclear how often LLMs can produce a valid proof when they find the correct answer.
 To investigate this, we first collect instances from the MathArena subset where models generate correct final answers, and then manually evaluate the validity of the accompanying proofs.
-This enables us to estimate the overall proof correctness rate P(correct proof)P($\text${correct proof}) and compare it with the final-answer accuracy P(correct final answer)P($\text${correct final answer}).
-In Fig.˜5, we report each model’s accuracy alongside the stricter metric requiring both a correct final answer and a valid proof.
-Despite Gemini-2.5-Pro, o4-mini, and o3 achieving similar final-answer accuracy, their proof correctness rates differ significantly.
+This enables us to estimate the overall proof correctness rate and compare it with the final-answer accuracy.
+Results
+In Fig.˜5, we report each model’s final-answer accuracy alongside the stricter metric requiring a valid proof.
+Despite Gemini-2.5-Pro, o4-mini, and o3 achieving similar final-answer accuracy, proof correctness differs significantly.
 Specifically, o3 performs notably worse, with only $59.5\%$ of its answers containing a correct proof.
-This substantial variation across models highlights that final-answer accuracy alone is not a reliable indicator of proof generation capability.
+This substantial difference between models shows that final-answer accuracy is not a reliable indicator of proof generation capability.
 5.5 Best-of-n Significantly Improves Performance
-Figure 6: Pass@n metric on a small subset of the OPC compared with some best-of-n selection strategies.
-Best-of-n selection—generating multiple outputs and selecting the best one—is a common strategy for improving LLM performance.
+((a)) Pass@n versus all other best-of-n methods.
+((b)) Selection of methods on a larger subset.
+Figure 6: Performance of the best-of-n selection strategies.
+Best-of-n selection strategies
+Best-of-n selection, generating multiple outputs and selecting the best one, is a common strategy to improve performance.
 We evaluate this approach using o4-mini by generating eight proofs per problem in the best-of-n subset and testing four selection methods:
-  • Discrete: o4-mini classifies each proof as correct or incorrect and selects any correct proof.
-  • Continuous: o4-mini scores proofs on a 0-77 scale and selects the one with the highest score.
-  • Rank (Bracket): A ranking method proposed by Liu et al.
-(2025).
-Proofs are judged pairwise by o4-mini in a knockout tournament until one proof remains.
-Note that Liu et al.
-(2025) finetuned their model on the judging task, while we use o4-mini itself.
+  • Discrete: o4-mini classifies each proof as correct or incorrect and selects any correct one.
+  • Continuous: o4-mini scores proofs on a 0-77 scale and selects one with the highest score.
+  • Rank (Bracket): A ranking method proposed by Liu et al. (2025). Proofs are judged pairwise by o4-mini in a knockout tournament until one proof remains.
   • Rank (Swiss): Inspired by Liu et al.
-(2025), proofs are ranked via pairwise comparisons in a Swiss round-robin tournament.
-Ratings are computed using the Bradley-Terry model (Bradley and Terry, 1952), and the answer with the highest rating is selected.
-See App.˜B for details of this method.
-Prompts for all methods can be found in App.˜E.
-Note that Rank (Bracket) only requires $O(n)$ comparisons, and is therefore as efficient as the discrete and continuous methods, while Rank (Swiss) requires $O(n^{2})$ comparisons, making it more expensive.
-In Fig.˜6, we compare the performance of these methods with the pass@n metric on the $60$ problems that contain human judgments for all eight proofs.
+(2025), proofs are paired in a Swiss round-robin tournament.
+Ratings are computed using the Bradley-Terry model (Bradley and Terry, 1952), and the proof with the highest rating is selected.
+See App.˜B for details.
+Prompts for all methods can be found in App.˜I.
+Complexity
+Note that Rank (Bracket) requires $O(n)$ comparisons, making it as efficient as the discrete method, while Rank (Swiss) requires $O(n^{2})$ comparisons, making it more expensive.
+Results on small subset
+In Fig.˜6(a), we compare the performance of these methods with the pass@n metric on the $60$ problems that contain human judgments for all eight proofs.
 We find that best-of-n selection strategies can significantly improve proof generation performance.
-Furthermore, the pairwise ranking methods significantly outperform the discrete and continuous methods by approximately $10\%$.
+Furthermore, the pairwise ranking methods outperform the discrete and continuous methods by approximately $10\%$.
 Notably, while discrete and continuous methods plateau after $n=5$, ranking approaches continue to scale.
 Finally, Rank (Swiss) slightly outperforms Rank (Bracket) by a $3\%$ margin.
-Figure 7: Best-of-n selection strategies on the entire best-of-n subset of the OPC.
-In Fig.˜7, we evaluate the performance of the best-of-n selection strategies on the entire best-of-n subset of the OPC, except for the Rank (Bracket) method, which was not evaluated on the full subset.
+Results on larger subset
+In Fig.˜6(b), we evaluate the performance of the best-of-n selection strategies on the $134$ problems of the best-of-n subset of the OPC [^1] A small bug in the Rank (Swiss) method caused incorrect selections for 18 questions.
+These are excluded from the analysis., except for the Rank (Bracket) method, which was not evaluated on the full subset.
 The improved performance of the ranking methods is confirmed on the entire best-of-n subset, with Rank (Swiss) improving accuracy by $17\%$.
+While the confidence intervals are relatively large due to the small number of problems in this problem set, all selection methods rely on the same underlying answers from o4-mini, making the relative performance differences significant.
+5.6 Potential Impact of Contamination
+Since many OPC problems are publicly available, there is a possibility that some were contaminated during training.
+In this section, we briefly explain why this plays an insignificant role in our results.
+Proof generation capabilities
+In App.˜C, we present a small experiment indicating that contamination has a smaller effect than problem difficulty.
+Nevertheless, it cannot be fully ruled out.
+In particular, the small gap in performance between Gemini-2.5-Pro and o4-mini in Section˜5.1 cannot be conclusively attributed to genuine performance differences.
+However, results on smaller, contamination-free datasets (Petrov et al., 2025; Balunović et al., 2025) also show Gemini-2.5-Pro outperforming o4-mini, supporting the interpretation that the performance gap is real.
+In all other sections, the conclusions are robust to contamination: the informal-formal gap is too large to be affected by small changes, the discrepancy between final-answer accuracy and proof correctness is based on problems from MathArena, a well-known benchmark with problems created in 2025, and best-of-n strategies are compared using the same model, making contamination irrelevant.
+Table 4: Model accuracy with and without providing the ground truth solutions. $\Delta$ shows the change in accuracy between the two settings.
+Judge pass@1 pass@1 (w/ Solution) $\Delta$
+GPT-5 89.3 89.0 -0.3
+Gemini-2.5-Pro 85.4 82.7 -2.7
+OPC-R1-8B 83.8 83.1 -0.7
+o4-mini 83.8 85.2 +1.4
+o3 83.1 85.9 +2.8
+Qwen3-235B-A22B 81.8 84.9 +3.1
+R1 80.9 80.8 -0.1
+R1-Qwen3-8B 70.7 75.4 +4.7
+Proof judging capabilities
+Data contamination poses a less significant risk for proof judging, since generated proofs cannot be present in the training data.
+The underlying problems and their official solutions, however, may have been seen during training, which could give models an advantage.
+To test this, we run a worst-case experiment where the ground-truth solution is provided alongside the proof to be judged.
+As shown in Table˜4, the resulting accuracy gains are small and non-significant, suggesting that access to the solution does little to improve error detection and that contamination has a limited impact on judging performance.
 6 Limitations
 While the OPC represents a significant advancement in the development and evaluation of LLM proof-generation capabilities, it is not without limitations.
-First, since most problems in the OPC are sourced from publicly available mathematical competitions, there is a risk of data contamination.
-Some of these problems may have been encountered during the training of the models.
-However, since the main purpose of the OPC is not related to benchmarking but rather to provide a large, human-validated dataset of LLM-generated proofs, this issue is less critical.
-Second, the majority of problems in the OPC are derived from high school level competitions.
-As a result, the dataset does not cover more advanced mathematical domains, such as undergraduate or research-level mathematics.
-7 Future Work
-The OPC offers a valuable resource for advancing research in proof generation with LLMs, supporting both the training and evaluation of models across a diverse set of mathematical problems.
-Our analysis also highlights several critical gaps and challenges that warrant further research.
-First, the significant disparity between formal and natural language proof generation—outlined in Section˜5.3—points to the need for more effective strategies to bridge this divide.
-Second, the OPC demonstrates that current benchmarks, which often rely solely on final answers, fail to capture the full complexity and quality of generated proofs, as discussed in Section˜5.4.
-This underscores the need for the development of a scalable benchmarking pipeline tailored to proof generation tasks.
-Finally, while our results show that best-of-n sampling strategies can meaningfully improve proof quality (see Section˜5.5), further research is required to better understand and optimize these methods.
+First, since dataset construction took place before the release of Grok-4 and GPT-5, these models are only included as judges.
+However, recent benchmarks suggest that these models perform similarly, perhaps slightly better, compared to Gemini-2.5-Pro (Balunović et al., 2025).
+Therefore, it does not affect the validity of our conclusions.
+Second, the majority of problems in the OPC are derived from high-school competitions.
+Thus, the dataset does not cover more advanced mathematical domains, such as research-level mathematics.
+We further outline potential directions for expanding the OPC in App.˜F.
 7 Conclusion
-In this work, we introduced the Open Proof Corpus (OPC), a large-scale, human-validated dataset of over 5,0005{,}000 LLM-generated proofs across more than 1,0001{,}000 problems.
-Designed to support training, evaluation, and benchmarking, the OPC provides a robust foundation for advancing research in automated proof generation.
-Using the OPC, we conducted a thorough evaluation of current LLM capabilities in mathematical reasoning.
-Our analysis reveals several key findings: (1) modern LLMs can generate a significant number of correct proofs; (2) some models approach human-level performance in judging proof correctness; (3) natural language proof generation substantially outperforms formal proof generation; (4) final-answer benchmarks are a poor proxy for proof quality; and (5) best-of-n sampling strategies significantly improve proof-generation accuracy.
-Using the OPC, we also trained an open-source model that achieves $88.1\%$ accuracy in judging proof correctness, matching the performance of the best closed-source model, Gemini-2.5-Pro.
+In this work, we introduced the Open Proof Corpus (OPC), a human-validated dataset comprising over 5,0005{,}000 LLM-generated proofs.
+Using the OPC, we addressed several open questions in the field, including (1) the gap between natural language and formal proof generation, (2) the relationship between final-answer accuracy and proof correctness, and (3) the effectiveness of best-of-n selection.
+We also trained a model that achieves $88.1\%$ accuracy in judging proof correctness, matching Gemini-2.5-Pro and approaching the performance of the top model, GPT-5.
+Reproducibility Statement
+We have included our dataset in the supplementary material, along with detailed descriptions of our methodology and experimental setup to ensure full reproducibility.
+In addition, we provide our code with step-by-step instructions in the supplementary material, enabling others to replicate our results and train the model from scratch.
+While the model itself is open-sourced, it is not included in the supplementary materials due to file size constraints.
+A small portion of the problems included in the dataset were obtained through private correspondence with the competition organizers (SMT 2025, $\approx 2\%$ of the data).
+They asked that we do not share these problems until they have released them publicly.
+We will therefore share these problems once they are made public by the competition organizers.
+Our results remain virtually unchanged if we exclude these problems from our dataset.
 Acknowledgements
@@
 The OPC contains over 1000 problems that were sourced from national and international competitions of varying difficulty.
-In Table˜4, we present the problem and sample distribution for each.
+In Table˜5, we present the problem and sample distribution for each.
 We also include the following additional information:
@@
   • Type: whether the competition is hosted internationally (I) or only nationally (N).
-  • Source: we link the source, from which we obtained the problems. Any source that is not publicly available was marked as "Private".
+  • Source: we link the source from which we obtained the problems. Any source that is not publicly available was marked as "Private".
   • Acc: the average accuracy of the best-performing model on the competition problems, which serves as a rough proxy for difficulty.
@@
 We fit the model to the outcomes of the round-robin tournament using maximum likelihood estimation, resulting in a rating for each proof. The proof with the highest rating is selected as the best.
-Appendix C Qualitative Observations
+Appendix C Additional Results
+In this section, we present further results, specifically addressing critical aspects of our evaluation: out-of-distribution performance of our trained model and the potential impact of data contamination.
+C.1 OPC-R1-8B performance on OOD problems
+During the training of OPC-R1-8B, our dataset primarily consisted of samples from high-school level competitions, aligning with the distribution of the test set.
+Crucially, undergraduate-level problems were explicitly excluded from training.
+To evaluate OPC-R1-8B’s generalization capabilities on out-of-distribution data, we evaluated its judging performance on 560 undergraduate-level solution samples sourced from the Putnam competition.
+Table 6: LLMs as proof graders on undergraduate-level problems.
+Cost for running the model on the entire subset is given in USD.
+Confidence intervals are 95% and computed using the large sample normal approximation.
+Judge pass@1 maj@5 Cost
+o4-mini $83.8\pm 3.1$ $85.1\pm 2.9$ 48.31
+GPT-5 $83.6\pm 3.1$ $84.0\pm 3.0$ 89.42
+R1 $78.3\pm 3.4$ $79.7\pm 3.3$ 51.86
+OPC-R1-8B $5.0\pm 3.6 77$$7.0\pm 3.5$ N/A
+R1-Qwen3-8B $72.1\pm 3.7$ $72.4\pm 3.7$ N/A
+Gemini-2.5-Pro $70.9\pm 3.8$ $76.7\pm 3.5$ 160.95
+As shown in Table˜6, OPC-R1-8B retains a notable improvement over its base model, despite undergraduate-level examples not being present in its testing set.
+While frontier models like o4-mini achieve higher performance, OPC-R1-8B is nevertheless competitive, performing on par with Gemini-2.5-Pro.
+This shows that the OPC data can be applied to mathematical domains beyond its scope.
+C.2 Problem Category and Difficulty Affect Incorrect Proof Rates
+We analyze the prevalence of incorrect proofs in the MathArena subset of the OPC.
+In particular, we investigate whether difficulty and problem type correlates with the likelihood of a model generating an incorrect proof.
+As shown in Fig.˜7, we find that both these factors significantly influence performance.
+In particular, higher difficulty levels correspond to a greater proportion of incorrect proofs, as illustrated in Fig.˜7(a).
+Additionally, among the four main problem categories, Geometry problems exhibit the highest rate of incorrect proofs.
+((a)) Problems are binned by difficulty level, with higher levels indicating more challenging problems.
+((b)) Problems are binned in the four main categories: Algebra, Combinatorics, Geometry, and Number Theory.
+Figure 7: Influence of difficulty level and problem category on the prevalence of incorrect proofs in MathArena.
+C.3 Data Contamination for Mathematical Proofs
+Test-set contamination is a prevalent challenge in evaluating the true reasoning capabilities of models, particularly in mathematics [Balunović et al., 2025, Zhang et al., 2024].
+Given that the vast majority of OPC problems are publicly sourced, we conduct one additional experiment to address potential contamination-related concerns.
+Competition-based breakdown
+First, we outline 2 categories of competitions: Standard and Non-standard.
+The former includes important competitions in the mathematics community, namely the IMO Shortlist and USAMO.
+Problems from these widely recognized competitions are more likely to appear in training sets, as they are widely reported on the internet and therefore represent the greatest potential risk for contamination.
+Figure 8: Average proof correctness on a sample of the OPC, split into Standard and Non-standard competitions.
+On the other hand, the Non-standard category consists of problems from the International Zhautykov Olympiad, the Swiss Mathematical Olympiad, and the Bulgarian seasonal competitions.
+A significant portion of these problems required translation from their original non-English statements, and they originate from less prevalent sources compared to the IMO and USAMO.
+Consequently, we anticipate a substantially lower prevalence of contamination within this problem set.
+As illustrated in Fig.˜8, model accuracy on Standard problems is notably lower than on their Non-standard counterparts.
+This observation indicates that problem difficulty is a considerably more dominant factor influencing model performance than the potential presence of test-set contamination.
+While this experiment alone cannot definitively discount the presence of test-set contamination within the OPC dataset for our Section˜5.1 results, it demonstrates that the models’ problem-solving capabilities extend significantly beyond memorization.
+This suggests that any contamination effect is likely small and not the primary driver of performance differences.
+C.4 Significance Analysis
+In Table˜7, we present the performance of various LLMs as proof judges, along with 95% confidence intervals using the large sample normal approximation.
+In Table˜8, we provide a detailed breakdown of judging accuracy by prover, again with 95% confidence intervals.
+These tables complement the main results presented in Section˜5.
+Table 7: LLMs as proof graders. Cost for running the model on the entire subset is given in USD. Confidence intervals are computed using the large sample normal approximation.
+Judge pass@1 maj@5 Cost
+Human $90.4\pm 3.4$ - N/A
+GPT-5 $89.3\pm 3.5$ $90.8\pm 3.3$ $117.77$
+Grok-4 $88.3\pm 3.7$ $89.8\pm 3.5$ $104.42$
+Gemini-2.5-Pro $5.4\pm 4.0 88$$8.1\pm 3.7$ $135.47$
+OPC-R1-8B $3.8\pm 4.2 88$$8.1\pm 3.7$ N/A
+o4-mini $83.8\pm 4.2$ $85.3\pm 4.1$ $29.57$
+o3 $83.1\pm 4.3$ $84.3\pm 4.2$ $93.3$
+Gemini-2.5-Flash $82.7\pm 4.3$ $86.0\pm 4.0$ $86.95$
+Qwen3-235B-A22B $81.8\pm 4.4$ $84.6\pm 4.1$ $3.79$
+R1 $80.9\pm 4.5$ $82.6\pm 4.3$ $27.70$
+R1-Qwen3-8B $70.7\pm 5.2$ $71.3\pm 5.2$ N/A
+Claude-4-Sonnet $70.6\pm 5.2$ $75.0\pm 5.0$ $28.21$
+Qwen3-8B $64.4\pm 5.5$ $63.6\pm 5.5$ N/A
+GPT-4.1 $61.4\pm 5.6$ $60.8\pm 5.6$ $20.33$
+Table 8: Judgement accuracy breakdown by prover with 95% confidence intervals.
+Prover Judge Gemini o4 o3 Qwen
+Gemini $79.4\pm 4.9$ $86.9\pm 4.1$ $85.9\pm 4.2$ $80.0\pm 4.9$
+o4 $87.1\pm 4.1$ $81.3\pm 4.8$ $84.8\pm 4.4$ $81.9\pm 4.7$
+o3 $91.6\pm 3.4$ $83.1\pm 4.6$ $76.9\pm 5.2$ $79.1\pm 5.0$
+Qwen $80.6\pm 4.9$ $84.1\pm 4.6$ $87.8\pm 4.1$ $84.4\pm 4.5$
+Appendix D Experimental Details
+In this section, we provide additional details regarding our experimental setup, including model training, hyperparameters, and evaluation protocols.
+D.1 Model Training
+Training data
+As outlined in Section˜5, we split the generic split of the OPC dataset into a train and test set.
+The training set consists of 1,733 proof samples, while the test set contains 293 proof samples.
+Importantly, we ensure that no problem statements overlap between the training and test sets, maintaining the integrity of our evaluation.
+Reinforcement learning
+We fine-tune R1-Qwen3-8B using GRPO [Shao et al., 2024] on the training set using the popular VERL framework [Sheng et al., 2024] with a learning rate of $10^{-6}$, a maximum response length of $14000$ tokens, $10$ rollouts per problem, and a batch size of $16$.
+We use the same prompt for training as for evaluation, as shown in Section˜I.3.
+More hyperparameters and training details can be found in the provided code.
+Importantly, we did not use the test set at any point during training or hyperparameter tuning, and only evaluated the final trained model once on the test set.
+D.2 Model Evaluation
+Hyperparameters and prompt
+For all models, we use the recommended hyperparameters outlined by the model providers.
+For open models, this usually involves using nucleus sampling with $p=0.95$ and a temperature of $0.6$.
+For closed models, we use the default settings provided by the API.
+Model-specific hyperparameters can be found in the provided code.
+Prompts used for evaluation are shown in App.˜I.
+Formal model performance
+To accurately compute the performance of Goedel-Prover-V2 [Lin et al., 2025b] on our subset of problems from Tsoukalas et al.
+[2024], we contacted the authors of the models to obtain the raw results for each problem.
+This allowed us to compute the exact accuracy on our selected problems, which was not reported in the original paper.
+Appendix E Qualitative Observations
 Building on our quantitative analysis, we present several qualitative insights drawn from our judges’ evaluations. These include common failure modes and notable model behaviors.
@@
 This results in incorrect final answers despite seemingly coherent reasoning.
-Below is an example of such a mistake made by DeepSeek-R1, where the incorrect sign reversal is marked with "**".
-[Uncaptioned image]  DeepSeek-R1 – Baltic Way 2004 P3
+Below is an example of such a mistake made by R1, where the incorrect sign reversal is marked with "**".
+[Uncaptioned image]  R1 – Baltic Way 2004 P3
 Struggles with synthetic geometry reasoning
@@
 [Uncaptioned image]  Grok 3 Mini – IMO SL 2000 G4
-Appendix D Grading Interface and Instructions
-This appendix outlines the grading interface and the accompanying instructions provided to judges.
-The full interface and documentation can be accessed online at https://judgeproofs.xyz/sample and https://judgeproofs.xyz/instructions.
+Appendix F Future Work
+The OPC offers a valuable resource for advancing research in proof generation with LLMs, supporting both the training and evaluation of models for this task.
+Our analysis also highlights several critical gaps and challenges that require further research.
+First, the significant disparity between formal and natural language proof generation shows the need for more effective strategies to bridge this gap.
+Second, the OPC shows that current benchmarks fail to capture the full complexity and quality of generated proofs as they often rely solely on final answers.
+Therefore, the development of a scalable benchmarking pipeline tailored to proof generation tasks is necessary.
+Finally, while our results show that best-of-n sampling strategies can meaningfully improve proof quality, further research is required to better understand and optimize these methods.
+Appendix G Statement for the Use of Large Language Models
+Beyond being the subject of our research, LLMs were used only as assistants to improve the clarity and quality of writing.
+They were not involved in aiding our research methodology, ideation, or for discovering related work.
+Appendix H Grading Interface and Instructions
+This appendix outlines the grading interface and the accompanying instructions provided to judges. The full interface and documentation can be accessed and reviewed in our supplementary material.
 Judge ID
-Each judge received a unique identifier, which served as their login credential on https://judgeproofs.xyz.
+Each judge received a unique identifier, which served as their login credential on our website.
 This ID was used to track grading progress while maintaining judge anonymity in the resulting dataset.
@@
 The grading interface was designed for clarity and ease of use.
-Figs.˜8, 9 and 10 illustrate its main components.
+Figs.˜9, 10 and 11 illustrate its main components.
 The left panel contains a navigation bar for switching between problems and competitions.
@@
 Figure 11: Grading form. Judges indicate correctness, provide a justification, highlight relevant content, and optionally express uncertainty or abstain.
-Appendix E Prompts
+Appendix I Prompts
 In this section, we provide the prompts used for various tasks in the OPC.
 The prompts are designed to be clear and concise, guiding the LLMs through the proof generation process while ensuring that they understand the requirements for correctness and clarity.
-In Section˜E.1, we present the prompts used for generating proofs.
-In Section˜E.2, we provide the prompt used to generate the LLM summary to aid human graders in identifying potential issues in the proof.
-In Section˜E.3, we present the prompt used for LLMs to judge the correctness of a proof, used in Section˜5.2.
-In Sections˜E.4, E.5 and E.6, we provide all prompts used for the LLMs in best-of-n sampling, as described in Section˜5.5.
-E.1 Proof Generation Prompt
+In Section˜I.1, we present the prompts used for generating proofs.
+In Section˜I.2, we provide the prompt used to generate the LLM summary to aid human graders in identifying potential issues in the proof.
+In Section˜I.3, we present the prompt used for LLMs to judge the correctness of a proof, used in Section˜5.2.
+In Sections˜I.5, I.6 and I.7, we provide all prompts used for the LLMs in best-of-n sampling, as described in Section˜5.5.
+I.1 Proof Generation Prompt
 The following prompt is used for problems with no final answer:
 Prompt The following prompt is used for problems with a final answer: Prompt
-E.2 Issues Interface Prompt
+I.2 Issues Interface Prompt
 Prompt
-E.3 LLM as Judge Prompt
+I.3 LLM as Judge Prompt
 Prompt
-E.4 LLM as Discrete Judge Prompt
+I.4 LLM as Judge Prompt with Ground Truth Solution
 Prompt
-E.5 LLM as Continuous Judge Prompt
+I.5 LLM as Discrete Judge Prompt
 Prompt
-E.6 LLM as Rank Judge Prompt
+I.6 LLM as Continuous Judge Prompt
 Prompt
+I.7 LLM as Rank Judge Prompt
+Prompt